# Chemistry Beyond Exact Solutions on a Quantum-Centric Supercomputer

### v1

@@
A universal quantum computer can be used as a simulator capable of predicting properties of diverse quantum systems [1]. Electronic structure problems in chemistry offer practical use cases around the hundred-qubit mark [2]. This appears promising since current quantum processors have reached these sizes. However, mapping these use cases onto quantum computers yields deep circuits, and for for pre-fault-tolerant quantum processors, the large number of measurements to estimate molecular energies leads to prohibitive runtimes [3]. As a result, realistic chemistry is out of reach of current quantum computers in isolation. A natural question is whether classical distributed computation can relieve quantum processors from parsing all but a core, intrinsically quantum component of a chemistry workflow. Here, we incorporate quantum computations of chemistry in a quantum-centric supercomputing architecture, using up to 6400 nodes of the supercomputer Fugaku to assist a Heron superconducting quantum processor. We simulate the N2 triple bond breaking in a correlation-consistent cc-pVDZ basis set, and the active-space electronic structure of [2Fe–2S] and [4Fe–4S] clusters [4], using 58, 45 and 77 qubits respectively, with quantum circuits of up to 10570 (3590 2-qubit) quantum gates. We obtain our results using a class of quantum circuits that approximates molecular eigenstates, and a hybrid estimator. The estimator processes quantum samples, produces upper bounds to the ground-state energy and wavefunctions supported on a polynomial number of states. This guarantees an unconditional quality metric for quantum advantage, certifiable by classical computers at polynomial cost. For current error rates, our results show that classical distributed computing coupled to quantum processors can produce good approximate solutions for practical problems beyond sizes amenable to exact diagonalization.
@@
Figure 1: Quantum-centric supercomputing architecture and workflow diagram. Left: We illustrate a simplified architecture used to execute our workflow. The architecture has a cluster with a quantum system alongside classical runtime nodes within an isolated environment. A workload management system controls hybrid quantum-classical jobs through middleware. Our workflow is distributed on a set of classical nodes. It includes standard quantum chemistry application routines such as computing electronic integrals, mapping to qubits, and preparing circuits to be executed. Right: Details of the classical post-processing step. The input is a set of noisy samples $\widetilde{\mathcal{X}}$ from the quantum execution that are processed with our configuration recovery step, using information from a vector ${\bf{n}}$ of reference orbital occupancies. The green inset shows an example where a configuration with $N_{\bf{x}}<N$ is corrected. The set of recovered configurations $\mathcal{X}_{\textrm{R}}$ is sub-sampled and distributed for projection and diagonalization on parallel classical nodes. A new average reference occupancy vector ${\bf{n}}$ is computed from the results, and the configuration recovery loop is repeated self-consistently until convergence.
@@ I Methods
for each spin-orbital tuple $(p\sigma)$, averaged on the $K$ batches. These occupancies are sent back to the configuration recovery step, and this entire self-consistent iteration is repeated until convergence. The initial guess for ${\bf{n}}$ used for the first round of recovery comes from the raw quantum samples in the correct particle sector. The configuration recovery routine can be seen effectively as a problem-informed clustering of a noisy signal around the occupations ${\bf{n}}$.
@@
To establish some insight on the configuration recovery, we analytically compute a lower bound on the probability of recovering electronic configurations that have been corrupted by noise, as a function of their Hamming distance from an integer rounding of ${\bf{n}}$, using a global depolarizing noise model ${\widetilde{P}_{\Psi}({\bf{x}})=\alpha P_{\Psi}({\bf{x}})+(1-\alpha)\frac{1}{2^{M}}}$, where $\alpha\in[0,1]$ is the parameter that controls the amount of noiseless quantum signal [29]. We find that the likelihood of recovery decays exponentially with the system size and also with the Hamming distance from the configuration that best approximates ${\bf{n}}$. Since classical CI methods also require exponential time to reach configurations with high Hamming distances from a reference state, it would be possible for the quantum-centric architecture presented here to perform a more efficient simulation at some finite system size and error rates.
@@ II Results
Figure 4: Experiments: Chemistry on large basis sets. (a) 58 qubits are used to model the N2 dissociation (cc-pVDZ basis set). (b) 45 qubits are used for the [2Fe-2S] cluster (TZP-DKH basis set) and (c) 77 qubits for the [4Fe-4S] cluster (TZP-DKH basis set). The top panels show a 3-dimensional representation of the geometry of each molecule. The middle panels show the qubits selected on a Heron quantum processor layout, following the same color convention as panel (b) in Fig. 3. The bottom panel in (a) shows the potential energy surface comparison, as well as the energy difference $\Delta E$ between the Heat-Bath Configuration Interaction (HCI) energy and the energies obtained from different methods, including the quantum estimator. The brown scatter plot shows the value of $E^{(k)}$ for all batches of configurations and the connected dots show $\min_{k}(E^{(k)})$. The bottom panel in (b) shows the energy-variance analysis for three different eigenstates that both HCI and our method find upon increasing the value of $d$, as labeled by the color bar. For each approximate eigenstate $|\psi^{(k)}\rangle$, the horizontal axis $\Delta H=\langle\psi^{(k)}|\hat{H}^{2}|\psi^{(k)}\rangle-\langle\psi^{(k)}|\hat{H}|\psi^{(k)}\rangle^{2}$. The bottom panel in (c) shows a comparison of the energy-variance analysis applied to quantum measurement outcomes and bitstrings (with the correct particle number) sampled from the uniform distribution. The DMRG energy in panels (b) and (c) is from Ref. [45].
@@ Acknowledgements.
We thank the IBM Quantum Service and Data team for help with the workflow execution. The authors acknowledge feedback and insightful conversations with Alberto Baiardi, Giuseppe Carleo, Garnet Kin-Lic Chan, Antonio Córcoles, Oliver Dial, Jay Gambetta, Artur Izmaylov, Yukio Kawashima, David Kremer, Isaac Lauer, Peter Love, Guglielmo Mazzola, Takahito Nakajima, Paul Nation, Hanhee Paik, Emily Pritchett, Max Rossmannek, Paul Schweigert, James E.T. Smith, Miles Stoudenmire, Andrew Wack, Christa Zoufal.
@@ II Methods
II.1 HPC quantum estimator
This section provides a detailed description of the HPC quantum estimator. We describe the subspace projection and diagonalization, the self-consistent configuration recovery scheme, a lower bound to the likelihood of recovering different families of electronic configurations, and different techniques to optimize circuit parameters to minimize the estimator energy.
@@ II.1.2 Self-consistent configuration recovery
We analyze the probability of recovering configurations using the configuration recovery scheme described above, as a function of their Hamming distance from ${\bf{x}}_{r}$. For this analysis, we consider a simple global depolarizing noise model, whose effect on the probability of sampled configurations is described by:
$\widetilde{P}_{\Psi}({\bf{x}})=\alpha P_{\Psi}({\bf{x}})+(1-\alpha)\frac{1}{2^{M}},$ (S16)
where $\alpha\in[0,1]$ is a parameter that quantifies the amount of quantum signal. While we use for simplicity a global depolarizing noise model, we foresee similar derivations could be done for more realistic noise models.
@@
$\begin{split}&P_{\textrm{recovery}}(M,N,b,\varepsilon)\geq\sum_{g=0}^{M-N-b}\sum_{h=0}^{N-b}\frac{\binom{M-N-b}{g}\binom{b}{h}}{2^{M}}\prod_{i=1}^{g}\left(\frac{i(1-\varepsilon)}{(1-\varepsilon)(b+i)+\varepsilon(N-b+h)}\right)\prod_{j=1}^{h}\left(\frac{j\varepsilon}{(1-\varepsilon)b+\varepsilon(N-b+j)}\right)\\ &+\sum_{g=0}^{N-b}\sum_{h=0}^{M-N-b}\frac{\binom{N-b}{g}\binom{b}{h}}{2^{M}}\prod_{i=1}^{g}\left(\frac{i(1-\varepsilon)}{(1-\varepsilon)(b+i)+\varepsilon(M-N-b+h)}\right)\prod_{j=1}^{h}\left(\frac{j\varepsilon}{(1-\varepsilon)b+\varepsilon(M-N-b+j)}\right).\end{split}$ (S23)
The two terms correspond to the two cases above, with the combinatorial fractions in each giving the probabilities of obtaining a configuration equivalent to the particular bitstring ${\bf{x}}_{\text{target}}$ plus $g$ random $0\rightarrow 1$ bit-flips in the 0-sector of ${\bf{x}}_{r}$ and $h$ random $0\rightarrow 1$ bit-flips in the 1-sector (first term), or $g$ random $1\rightarrow 0$ bit-flips in the one-sector of the HF state and $h$ random $1\rightarrow 0$ bit-flips in the zero-sector (second term).
Fig. S2 shows the lower bound in Eq. (S23) to the recovery probability, for $\varepsilon=0.1$ and $N=10$, as a function of the number of qubits $M$. We observe that the lower bound decays as an exponential of the number of qubits and that the recovery of configurations with small values of $b$ is more likely than the recovery of configurations defined by large $b$ values. We remark that the configuration recovery probability for configurations close to ${\bf{x}}_{r}$ is substantially larger than the probability of obtaining the same samples by sampling from the uniform distribution, which would be the limit of no quantum signal. We argue that this feature helps in improving molecular energies over using simple post-selection over correct particle sector.
Figure S2: Lower bound to the probability of recovering a particular bitstring with Hamming distance $2b$ from a reference approximation ${\bf{x}}_{r}$ bistring from a bistring sampled from the uniform distribution. The lower bound is shown as a function of the number of qubits. Different values of $b$ are shown, as indicated by the colorbar. The dashed black curve shows $1/2^{M}$, which is the probability of obtaining the desired bitsting directly from the uniform distribution. The inset shows the lower bound as a function of $b$ for $M=52$ qubits.
@@ II.1.4 Orbital optimization
See Sec. IV.3 to see the effect of orbital optimizations in the accuracy of the HPC quantum estimator (with configuration recovery). The section shows results run on measurement outcomes obtained from the Heron device, to study the dissociation of N2 at cc-pVDZ level of theory.
@@ II.2.1 Sample complexity
${\langle{G_{m}}|}\hat{H}{|{G_{m}}\rangle}$ $\leq{\langle{G}|}\hat{H}{|{G}\rangle}+2\|\hat{H}\|\left\|{|{G}\rangle}-{|{G_{m}}\rangle}\right\|^{2}$
$\leq{\langle{G}|}\hat{H}{|{G}\rangle}+4\|\hat{H}\|(1-\sqrt{\alpha_{m}}).$ (S47)
@@
${\langle{\tilde{G}}|}\tilde{H}{|{\tilde{G}}\rangle}-{\langle{G}|}\hat{H}{|{G}\rangle}\leq 4\|\hat{H}\|(1-\sqrt{\alpha_{m}}).$ (S49)
@@
${\langle{\tilde{G}}|}\tilde{H}{|{\tilde{G}}\rangle}-{\langle{G}|}\hat{H}{|{G}\rangle}\leq 4\|\hat{H}\|(1-\sqrt{\alpha_{m}})$ (S51)
with probability at least $1-\eta$. This result allows us to estimate the sufficient number of samples $N_{s}$ given a distribution of the configurations in the ground state. For example, if $P_{I}\propto e^{-I}$ decays exponentially, we have $1-\sqrt{\alpha_{m}}\propto e^{-m}$ and $\beta_{m}\propto e^{-m}$. At $m\approx c\log M$, where $c$ is a parameter to be determined, choosing $N_{s}=\Omega(M^{c})$ is sufficient to guarantee an energy error at most $\mathcal{O}\left({\|\hat{H}\|/M^{c}}\right)$. We have the freedom to choose $c$ large enough so that $\|\hat{H}\|/M^{c}\ll 1$. In particular, if $\|\hat{H}\|=\mathcal{O}\left({M^{4}}\right)$, we would choose $c>4$. In short, the sufficient number of samples $N_{s}$ scales only polynomially with the number of qubits $M$.
This favorable scaling persists even if $P_{I}\propto 1/I^{\gamma}$ decays algebraically for some constant $\gamma>1$. In this case, we have $1-\sqrt{\alpha_{m}}\propto 1/m^{\gamma-1}$ and $\beta_{m}\propto 1/m^{\gamma}$. Therefore, choosing $N_{s}=\Omega(m^{\gamma-1})$, we can guarantee the error is at most $O\big{(}{\|\hat{H}\|/m^{\gamma}}\big{)}$. Again, we have the freedom to choose $m\propto M^{c}$ for some $c$ large enough so that $\|\hat{H}\|/M^{c\gamma}\ll 1$. With this choice of $m$, the number of samples only needs to scale polynomially with $M$ as $O\left(M^{c(\gamma-1)}\right)$.

### v2

@@
A universal quantum computer can be used as a simulator capable of predicting properties of diverse quantum systems [1]. Electronic structure problems in chemistry offer practical use cases around the hundred-qubit mark [2]. This appears promising since current quantum processors have reached these sizes. However, mapping these use cases onto quantum computers yields deep circuits, and for pre-fault-tolerant quantum processors, the large number of measurements to estimate molecular energies leads to prohibitive runtimes [3]. As a result, realistic chemistry is out of reach of current quantum computers in isolation. A natural question is whether classical distributed computation can relieve quantum processors from parsing all but a core, intrinsically quantum component of a chemistry workflow. Here, we incorporate quantum computations of chemistry in a quantum-centric supercomputing architecture, using up to 6400 nodes of the supercomputer Fugaku to assist a quantum computer with a Heron superconducting processor. We simulate the N2 triple bond breaking in a correlation-consistent cc-pVDZ basis set, and the active-space electronic structure of [2Fe–2S] and [4Fe–4S] clusters [4], using 58, 45 and 77 qubits respectively, with quantum circuits of up to 10570 (3590 2-qubit) quantum gates. We obtain our results using a class of quantum circuits that approximates molecular eigenstates, and a hybrid estimator. The estimator processes quantum samples, produces upper bounds to the ground-state energy and wavefunctions supported on a polynomial number of states. This guarantees an unconditional quality metric for quantum advantage, certifiable by classical computers at polynomial cost. For current error rates, our results show that classical distributed computing coupled to quantum computers can produce good approximate solutions for practical problems beyond sizes amenable to exact diagonalization.
@@
Figure 1: Quantum-centric supercomputing architecture and sample-based quantum diagonalization workflow diagram. Left: We illustrate a simplified architecture used to execute our workflow. The architecture has a cluster with a quantum system alongside classical runtime nodes within an isolated environment. A workload management system controls hybrid quantum-classical jobs through middleware. Our workflow is distributed on a set of classical nodes. It includes standard quantum chemistry application routines such as computing electronic integrals, mapping to qubits, and preparing circuits to be executed. Right: Details of the classical post-processing step. The input is a set of noisy samples $\widetilde{\mathcal{X}}$ from the quantum execution that are processed with our configuration recovery step, using information from a vector ${\bf{n}}$ of reference orbital occupancies. The green inset shows an example where a configuration with $N_{\bf{x}}<N$ is corrected. The set of recovered configurations $\mathcal{X}_{\textrm{R}}$ is sub-sampled and distributed for projection and diagonalization on parallel classical nodes. A new average reference occupancy vector ${\bf{n}}$ is computed from the results, and the configuration recovery loop is repeated self-consistently until convergence.
@@ I Methods
for each spin-orbital tuple $(p\sigma)$, averaged on the $K$ batches. These occupancies are sent back to the configuration recovery step, and this entire self-consistent iteration is repeated until convergence, realizing a sample-based quantum diagonalization (SQD) of the target Hamiltonian. The initial guess for ${\bf{n}}$ used for the first round of recovery comes from the raw quantum samples in the correct particle sector. The configuration recovery routine can be seen effectively as a problem-informed clustering of a noisy signal around the occupations ${\bf{n}}$.
@@
To establish some insight on the SQD workflow in relation to classical CI methods, consider that any CI method needs to identify useful configurations in an efficient way. This motivates the search for quantum circuits that can sample more effectively than classical methods. In the Supplementary Information [29] we show that at least one example of such circuits exists. We produce a circuit that searches the space of electronic configurations more efficiently than some flavors of SCI on an strongly correlated iron-sulfur cluster described by a (30 electron, 20 orbital) active space.
@@ II Results
Figure 4: Experiments: Chemistry on large basis sets. (a) 58 qubits are used to model the N2 dissociation (cc-pVDZ basis set). (b) 45 qubits are used for the [2Fe-2S] cluster (TZP-DKH basis set) and (c) 77 qubits for the [4Fe-4S] cluster (TZP-DKH basis set). The top panels show a 3-dimensional representation of the geometry of each molecule. The middle panels show the qubits selected on a Heron quantum processor layout, following the same color convention as panel (b) in Fig. 3. The bottom panel in (a) shows the potential energy surface comparison, as well as the energy difference $\Delta E$ between the Heat-Bath Configuration Interaction (HCI) energy and the energies obtained from different methods. The brown scatter plot shows the value of $E^{(k)}$ for all batches of configurations and the connected dots show $\min_{k}(E^{(k)})$. The bottom panel in (b) shows the energy-variance analysis for three different eigenstates that both HCI and our method find upon increasing the value of $d$, as labeled by the color bar. For each approximate eigenstate $|\psi^{(k)}\rangle$, the horizontal axis $\Delta H=\langle\psi^{(k)}|\hat{H}^{2}|\psi^{(k)}\rangle-\langle\psi^{(k)}|\hat{H}|\psi^{(k)}\rangle^{2}$. The bottom panel in (c) shows a comparison of the energy-variance analysis applied to quantum measurement outcomes and bitstrings (with the correct particle number) sampled from the uniform distribution. The DMRG energy in panels (b) and (c) is from Ref. [45].
@@ Acknowledgements.
We thank the IBM Quantum Service and Data team for help with the workflow execution. The authors acknowledge feedback and insightful conversations with Alberto Baiardi, Sergey Bravyi, Giuseppe Carleo, Garnet Kin-Lic Chan, Antonio Córcoles, Oliver Dial, Jay Gambetta, Artur Izmaylov, Caleb Johnson, Yukio Kawashima, David Kremer, Isaac Lauer, Peter Love, Guglielmo Mazzola, Takahito Nakajima, Paul Nation, Hanhee Paik, Emily Pritchett, Max Rossmannek, Paul Schweigert, James E.T. Smith, Miles Stoudenmire, Nobuyuki Yoshioka, Andrew Wack, and Christa Zoufal.
Data availability
All data supporting the findings are available from the corresponding authors upon reasonable request.
@@ II Methods
II.1 Sample-based Quantum Diagonalization
This section provides a detailed description of the Sample-based quantum diagonalization (SQD) procedure. We describe the subspace projection and diagonalization, the self-consistent configuration recovery scheme, a lower bound to the likelihood of recovering different families of electronic configurations, and different techniques to optimize circuit parameters to minimize the estimator energy.
@@ II.1.2 Self-consistent configuration recovery
Let $\mathcal{U}=\mathcal{U}_{L}\cdots\mathcal{U}_{2}\cdot\mathcal{U}_{1}$ denote the ideal unitary channel that produces $\psi=\mathcal{U}(|0\rangle\langle 0|)$. Here, $\mathcal{U}_{i}$ denotes a two qubit unitary. Let us analyze the effect of noise on $\mathcal{U}$. Suppose that each two-qubit gate $\mathcal{U}_{i}$ is followed by a Pauli noise channel that preserves the state with probability $p$, i.e, $\mathcal{P}(\rho)=p\rho+(1-p)P\rho P$ with $P$ denoting a Pauli operator. Then after $L$ quantum operations, the noisy output state $\widetilde{\rho}$ is given by
$\widetilde{\rho}=p^{L}\rho+\sigma\leavevmode\nobreak\ ,$ (S16)
where $\sigma$ denotes the component of the state affected from non-identity Pauli paths.
From Eq. (S16), it follows that the number of experiments needed to sample from the ideal state $\psi$ scales as $1/p^{L}$. Let $p=e^{-\lambda}$, where $\lambda$ denotes the noise rate. Then the number of shots needed for sampling bitstrings from $\psi$ scales as $e^{\lambda L}$, which is exponential in the number of noisy operations and the noise rate. In general, without performing quantum error correction, any quantum algorithm employed on noisy devices is limited to a maximum number of operations $L$ that scales inversely proportional to the noise rate $\lambda$. Note that for our application, one might also get correct bitstrings from $\sigma$, as defined in Eq. (S16), which can decrease the number of experiments needed to obtain good bitstrings.
Above we discussed the effect of noise in obtaining samples from the ideal state. We now analyze the probability of recovering configurations using the configuration recovery scheme described earlier, as a function of their Hamming distance from ${\bf{x}}_{r}$. In general, this analysis depends on the noise model considered. To simplify the analysis, we consider a simple global depolarizing noise model, whose effect on the probability of sampled configurations is described by:
$\widetilde{P}_{\Psi}({\bf{x}})=\alpha P_{\Psi}({\bf{x}})+(1-\alpha)\frac{\mathbb{I}}{2^{M}},$ (S17)
where $\alpha\in[0,1]$ is a parameter that quantifies the amount of quantum signal. Note that at $\alpha=1$, we exactly sample from the ideal state and, therefore, we do not need to perform the recovery process. While we use for simplicity a global depolarizing noise model, we expect that the analysis can be generalized to account for more realistic noise models, including the one defined in Eq. (S16).
@@
$\begin{split}&P_{\textrm{recovery}}(M,N,b,\varepsilon)\geq\sum_{g=0}^{M-N-b}\sum_{h=0}^{N-b}\frac{\binom{M-N-b}{g}\binom{b}{h}}{2^{M}}\prod_{i=1}^{g}\left(\frac{i(1-\varepsilon)}{(1-\varepsilon)(b+i)+\varepsilon(N-b+h)}\right)\prod_{j=1}^{h}\left(\frac{j\varepsilon}{(1-\varepsilon)b+\varepsilon(N-b+j)}\right)\\ &+\sum_{g=0}^{N-b}\sum_{h=0}^{M-N-b}\frac{\binom{N-b}{g}\binom{b}{h}}{2^{M}}\prod_{i=1}^{g}\left(\frac{i(1-\varepsilon)}{(1-\varepsilon)(b+i)+\varepsilon(M-N-b+h)}\right)\prod_{j=1}^{h}\left(\frac{j\varepsilon}{(1-\varepsilon)b+\varepsilon(M-N-b+j)}\right)\equiv F(M,N,b,\epsilon).\end{split}$ (S24)
The two terms correspond to the two cases above, with the combinatorial fractions in each giving the probabilities of obtaining a configuration equivalent to the particular bitstring ${\bf{x}}_{\text{target}}$ plus $g$ random $0\rightarrow 1$ bit-flips in the 0-sector of ${\bf{x}}_{r}$ and $h$ random $0\rightarrow 1$ bit-flips in the 1-sector (first term), or $g$ random $1\rightarrow 0$ bit-flips in the one-sector of the HF state and $h$ random $1\rightarrow 0$ bit-flips in the zero-sector (second term). In summary, the probability of configuration recovery to output a configuration without error is lower bounded by $\alpha+(1-\alpha)F(M,N,b,\epsilon)$, where $F(M,N,b,\epsilon)$ is defined in Eq. (S24). Note that this lower bound approaches one as $\alpha\rightarrow 1$.
Fig. S2 shows the lower bound in Eq. (S24) to the recovery probability, for $\varepsilon=0.1$ and $N=10$, as a function of the number of qubits $M$. We observe that the lower bound decays as an exponential of the number of qubits and that the recovery of configurations with small values of $b$ is more likely than the recovery of configurations defined by large $b$ values. We remark that the configuration recovery probability for configurations close to ${\bf{x}}_{r}$ is substantially larger than the probability of obtaining the same samples by sampling from the uniform distribution, which would be the limit of no quantum signal. We argue that this feature helps in improving molecular energies over using simple post-selection over correct particle sector. Additionally, Fig. S2 shows the fraction of configurations obtained from the configuration recovery procedure for different values of $b$ on actual hardware data on the N2 molecule on different bases sets and correspondingly different qubit numbers. The fractions are orders of magnitude larger than the lower bound and do not appear to have a strong dependence on the system size. This observation highlights that the lower bound is quite loose and in practice the performance is better.
Figure S2: Lower bound to the probability of recovering a particular bitstring with Hamming distance $2b$ from a reference approximation ${\bf{x}}_{r}$ bistring from a bistring sampled from the uniform distribution. The lower bound is shown as a function of the number of qubits. Different values of $b$ are shown, as indicated by the colorbar. The dashed black curve shows $1/2^{M}$, which is the probability of obtaining the desired bitsting directly from the uniform distribution. The circles connected by lines show the fraction of configurations for different families of bitstrings (as indicated by the colorbar) obtained by running the configuration recovery procedure on the samples from the quantum device on the N2 molecule for the 6-31-G and cc-pVDZ bases sets. The inset shows the lower bound as a function of $b$ for $M=52$ qubits.
@@ II.1.4 Orbital optimization
See Sec. V.3 to see the effect of orbital optimizations in the accuracy of the HPC quantum estimator (with configuration recovery). The section shows results run on measurement outcomes obtained from the Heron device, to study the dissociation of N2 at cc-pVDZ level of theory.
@@ II.2.1 Sample complexity
${\langle{G_{m}}|}\hat{H}{|{G_{m}}\rangle}$ $\leq{\langle{G}|}\hat{H}{|{G}\rangle}+2\|\hat{H}\|\left\|{|{G}\rangle}-{|{G_{m}}\rangle}\right\|$
$\leq{\langle{G}|}\hat{H}{|{G}\rangle}+2\sqrt{2}\|\hat{H}\|(1-\sqrt{\alpha_{m}})^{1/2}.$ (S48)
@@
${\langle{\tilde{G}}|}\tilde{H}{|{\tilde{G}}\rangle}-{\langle{G}|}\hat{H}{|{G}\rangle}\leq 2\sqrt{2}\|\hat{H}\|(1-\sqrt{\alpha_{m}})^{1/2}.$ (S50)
@@
${\langle{\tilde{G}}|}\tilde{H}{|{\tilde{G}}\rangle}-{\langle{G}|}\hat{H}{|{G}\rangle}\leq 2\sqrt{2}\|\hat{H}\|(1-\sqrt{\alpha_{m}})^{1/2}$ (S52)
with probability at least $1-\eta$. This result allows us to estimate the sufficient number of samples $N_{s}$ given a distribution of the configurations in the ground state. For example, if $P_{I}\propto e^{-I}$ decays exponentially, we have $1-\sqrt{\alpha_{m}}\propto e^{-m}$ and $\beta_{m}\propto e^{-m}$. At $m\approx 2c\log M$, where $c$ is a parameter to be determined, choosing $N_{s}=\Omega(M^{2c})$ is sufficient to guarantee an energy error at most $\mathcal{O}\left({\|\hat{H}\|/M^{c}}\right)$. We have the freedom to choose $c$ large enough so that $\|\hat{H}\|/M^{c}\ll 1$. In particular, if $\|\hat{H}\|=\mathcal{O}\left({M^{4}}\right)$, we would choose $c>4$. In short, the sufficient number of samples $N_{s}$ scales only polynomially with the number of qubits $M$.
This favorable scaling persists even if $P_{I}\propto 1/I^{\gamma}$ decays algebraically for some constant $\gamma>1$. In this case, we have $1-\sqrt{\alpha_{m}}\propto 1/m^{\gamma-1}$ and $\beta_{m}\propto 1/m^{\gamma}$. Therefore, choosing $N_{s}=\Omega(m^{\gamma})$, we can guarantee the error is at most $O\big{(}{\|\hat{H}\|/m^{(\gamma-1)/2}}\big{)}$. Again, we have the freedom to choose $m\propto M^{c}$ for $c$ large enough so that $\|\hat{H}\|/M^{c(\gamma-1)/2}\ll 1$. With this choice of $m$, the number of samples only needs to scale polynomially with $M$ as $O\left(M^{c\gamma}\right)$.